Inverse Functions.
1)Explain how to find the inverse of a function when the function is given in each of the following formats.
a)As a table of inputs and outputs.
b)As a graph.
c)As a equation.


2)Explain how to determine if two functions in equation form are inverses of each other and give a sample pair of functions for your classmates to practice determining if they are functions.

Answer by Theo(13342)

let's take a look at how we derive inverse functions by equation.

if y = 5*x, then you find the inverse qeuation by replacing y with x and x with y and solving for y.

y = 5*x becomes x = 5*y

solve for y to get y = x/5.

the inverse equation of y = 5x is y = x/5.

that's by equation.

now by table.

take the same equation of y = 5x and make a table.

you get:

(x,y) = (1,5), (2,10), (3,15), (4,20), etc.

now replace x with y and replace y with x to get:

(x,y) = (5,1), (10,2), (15,3), (20,4), etc.

the first table gets you the equation of y = 5x.

the second table gets you the equation of y = x/5.

all you do is interchange the x valuea and the y values and you have a table for the inverse equation.

graphically, you would plot the (x,y) points for y = 5x.

for each (x,y) point, you would plot the corresponding point of (y,x).

if you then draw the line y = x on your graph you will see that the points of (y,x) are reflections of the points (x,y) about the line y = x.

here's your graph.

you find the reflected points about the line y = x by drawing a line vertical to that line (represented by y = -x) and then finding the intersection of that vertical line with the reflected lines of y = 5x and y = x/5.

the point of intersection will be interchanged.

(10,2) becomes (2,10) and (-10,-2) becomes (-2,-10) as seen on the graph.


$$$